A credit company randomly selected 50 contested items and recorded the dollar amount being contested. These contested items had sample mean ?ത = 95.74 dollars and s = 24.63 dollars. Construct a point estimate for the population mean contested amount, μ, and give its 90% error margin
Answer #1

The point estimate for population mean is = 95.74

DF = 50 - 1 = 49

At 90% confidence level, the critical value is t* = 1.677

Margin of error = t* * s/

= 1.677 * 24.63/

= 5.8413